# CogniBench: A Legal-inspired Framework and Dataset for Assessing Cognitive Faithfulness of Large Language Models

### v2

@@ Abstract
Faithfulness hallucination are claims generated by a Large Language Model (LLM) not supported by contexts provided to the LLM. Lacking assessment standard, existing benchmarks only contain “factual statements” that rephrase source materials without marking “cognitive statements” that make inference from the given context, making the consistency evaluation and optimization of cognitive statements difficult. Inspired by how an evidence is assessed in the legislative domain, we design a rigorous framework to assess different levels of faithfulness of cognitive statements and create a benchmark dataset where we reveal insightful statistics. We design an annotation pipeline to create larger benchmarks for different LLMs automatically, and the resulting larger-scale CogniBench-L dataset can be used to train accurate cognitive hallucination detection model. We release our model and dataset at: https://github.com/FUTUREEEEEE/CogniBench
@@ 1 Introduction
However, as LLMs evolve, LLMs increasingly generate cognitive statements— such as inferences, evaluations, and explanations that extend beyond the verbatim context Webb et al. (2023); Zheng et al. (2023); Xu et al. (2023). This is essential for increasing applications in medicine Singhal et al. (2025), legislation Singhal et al. (2025), and finance Yao et al. (2024). For instance, a medical AI must not only recall symptoms from a patient report but also draw diagnostic conclusions. Put in a broader perspective, we categorize LLM-generated statements based on Bloom’s Taxonomy Bloom et al. (1956): factual statements (recalling context) and cognitive statements (applying, explaining, or evaluating knowledge), as exemplified in Fig. 1. Existing benchmarks Mishra et al. (2024); Niu et al. (2023); Hasan et al. (2021), which prioritize factual consistency, are primarily dominated by “factual statement” as shown in Fig. 2.
@@ A.12 Compare between CogniBench and existing benchmarks
We compare the query context and response from RAGTruth Niu et al. (2023) and CNN/Daily Mail See et al. (2017) in Fig. 14, Both RAGTruth and CNN/Daily Mail dataset are composed maily with factual statement RAGTruth answers focuses more on rephrase the information from the retrieved passages. CNN/Daily Mail excels in news summarization, where the response retrieval key factual information form context to generate a condensing article.

### v3

@@ Abstract
Faithfulness hallucinations are claims generated by a Large Language Model (LLM) not supported by contexts provided to the LLM. Lacking assessment standards, existing benchmarks focus on “factual statements” that rephrase source materials while overlooking “cognitive statements” that involve making inferences from the given context. Consequently, evaluating and detecting the hallucination of cognitive statements remains challenging. Inspired by how evidence is assessed in the legal domain, we design a rigorous framework to assess different levels of faithfulness of cognitive statements and introduce the CogniBench dataset where we reveal insightful statistics. To keep pace with rapidly evolving LLMs, we further develop an automatic annotation pipeline that scales easily across different models. This results in a large-scale CogniBench-L dataset, which facilitates training accurate detectors for both factual and cognitive hallucinations. We release our model and datasets at: https://github.com/FUTUREEEEEE/CogniBench
@@ 1 Introduction
However, as LLMs evolve, LLMs increasingly generate cognitive statements— such as inferences, evaluations, and explanations that extend beyond the verbatim context Webb et al. (2023); Zheng et al. (2023); Xu et al. (2023). This is essential for increasing applications in medicine Singhal et al. (2025), legislation Singhal et al. (2025), and finance Yao et al. (2024). For instance, a medical AI must not only recall symptoms from a patient report but also draw diagnostic conclusions. Put in a broader perspective, we categorize LLM-generated statements based on Bloom’s Taxonomy Bloom et al. (1956): factual statements (restate facts from context) and cognitive statements (applying, explaining, or evaluating knowledge), as exemplified in Fig. 1. Existing benchmarks Mishra et al. (2024); Niu et al. (2023); Hasan et al. (2021), which prioritize factual consistency, are primarily dominated by “factual statement” as shown in Fig. 2.
@@ A.12 Compare between CogniBench and existing benchmarks
We compare the query context and response from RAGTruth Niu et al. (2023) and CNN/Daily Mail See et al. (2017) in Fig. 14, Both RAGTruth and CNN/Daily Mail dataset are composed mainly with factual statement RAGTruth answers focuses more on rephrase the information from the retrieved passages. CNN/Daily Mail excels in news summarization, where the response retrieval key factual information form context to generate a condensing article.